OpenStudy (anonymous):

what is the method for 7over 2 plus 3 over 4

OpenStudy (ccswims):

\[\frac{ 7 }{ 2 } + \frac{ 3 }{ 4 }\]Now you must multiply 2 by 2 to go into 4. Whatever you do to the denominator you must do to the numerator, so \[\frac{ 14 }{ 4 } + \frac{ 3 }{ 4 }= \frac{ 17 }{ 4 }\]

OpenStudy (ccswims):

Simplify to \[4\frac{ 1 }{ 4 }\]